OpenStudy (anonymous):

A block of ice loses one third of its weight every hour. After 1 hour, the block weighs 60 pounds. How much did the block weigh to begin with? A. 90 lbs B. 120 lbs C. 100 lbs

OpenStudy (anonymous):

Hence, if the original weight is "W", then (2/3)W = 60 Solving for "W", W = 60 * 3/2 W = 90 lbs.